# Orbital classification in rotating bar potentials using an empirical proxy of the second integral of motion

### v1

@@ I Introduction
Our main motivation is to investigate the effectiveness of $\mathrm{CAM}$ for 2D orbital classification in rotating bar potentials. While QS21 focused primarily on the empirical and numerical properties of $\mathrm{CAM}$, we begin by investigating its analytical implications within the Freeman bar model, aiming to gain insight into the physical meaning of this proxy. Next, we evaluate the performance of $\mathrm{CAM}$ for orbital classification across three representative rotating bar potentials: the Freeman bar, the logarithmic bar, and the Ferrers bar. In addition to the $\overline{L_{z}}-\sigma_{L_{z}}$ plane studied by QS21, we introduce a new parameter plane defined by $\mathrm{CAM}$ and the root-mean-square radius (${R_{\mathrm{rms}}}=\sqrt{\overline{R^{2}}}$) [^1] Throughout this paper, we use $R_{\mathrm{rms}}$} to characterize the orbital size, but we do confirm that the mean absolute radius ($\overline{|R|}$) produces comparable results.. We first explore the behavior of iso-$E_{\mathrm{J}}$ and periodic orbits within the $\mathrm{CAM}$ framework, finding that periodic orbits form well-defined branches in the $\mathrm{CAM}-{R_{\mathrm{rms}}}$ plane that effectively partition the space into domains corresponding to specific orbital families. We then extend the analysis to test-particle orbits rather than orbits constrained by fixed Jacobi energies, thereby demonstrating the applicability of this method to realistic $N$-body simulations. Our analyses show that different orbital families exhibit a well-ordered distribution in the $\mathrm{CAM}-{R_{\mathrm{rms}}}$ plane. These results highlight CAM as a robust and efficient tool for orbital classification.
@@ III Iso-$E_{\mathrm{J}}$ orbits and periodic orbits in the $\mathrm{CAM}$ framework
In the Freeman bar potential, the iso-$E_{\mathrm{J}}$ orbits in the $\overline{L_{z}}-\sigma_{L_{z}}$ plane are shown in the black curves in the left panel of Figure 1, forming the upper portion of an ellipse. The periodic $x_{1}$ and $x_{4}$ orbits in this model, however, exhibit zero angular momentum dispersion and constant $L_{z}$ as $X_{\alpha}$ or $X_{\beta}$ equals to 0 (see Equation 9), causing their $\mathrm{CAM}$ values to diverge. As a result, their properties cannot be meaningfully analyzed within the $\mathrm{CAM}$ framework. For this reason, we also consider two additional representative rotating bar potentials: the logarithmic bar and the Ferrers bar in J. A. Sellwood & A. Wilkinson (1993). The logarithmic bar model admits only the $x_{1}$ and $x_{4}$ orbital families, whereas the Ferrers bar potential also supports the $x_{2}$ orbital family. Orbits in these two potentials are integrated for a timespan of around $10$ dynamical timescales using the python package Agama (E. Vasiliev, 2018). The periodic orbits are obtained by launching particles along the $x$-axis at fixed $E_{\mathrm{J}}$ and identifying those that return to their initial phase-space coordinates.
@@
In addition to the $\overline{L_{z}}-\sigma_{L_{z}}$ plane examined in the original CAM study QS21, we also explore alternative parameter spaces that directly incorporate the $\mathrm{CAM}$ value. Since CAM serves as a proxy for $I_{2}$ in addition to $E_{\mathrm{J}}$, the $\mathrm{CAM}-E_{\mathrm{J}}$ plane, resembling the integral-of-motion space, naturally provides a framework for orbital classification. However, because orbits inside and outside the corotation radius overlap in the $\mathrm{CAM}-E_{\mathrm{J}}$ plane, an issue discussed in detail in Section IV.2 later, we instead adopt the $\mathrm{CAM}-{R_{\mathrm{rms}}}$ plane. We find that this parameter space provides a clear separation between orbital families and serves as a critical diagnostic for orbital classification within the $\mathrm{CAM}$ framework.
@@ III.1 $x_{1}$ and $x_{4}$ orbits
$\Phi(x,y)=\frac{1}{2}\ln(R_{c}^{2}+x^{2}+\frac{y^{2}}{b^{2}}).$ (22)
We take $R_{c}=0.1$ and $b=0.84$, with a pattern speed of $\Omega_{\mathrm{b}}=1$. This model only supports $x_{1}$ and $x_{4}$ orbital families, while $x_{2}$ and $x_{3}$ orbital families are absent.
@@
Figure 2: Orbital distributions of the logarithmic bar potential, shown in the ${\overline{L}_{z}}-\sigma_{L_{z}}$ plane (left) and the $\mathrm{CAM}-{R_{\mathrm{rms}}}$ plane (right), color-coded by $E_{\mathrm{J}}$. Representative periodic orbits, displayed with their morphological shapes, are superimposed on the distribution, while quasi-periodic orbits appear as colored curves composed of discrete points. Distinct branches corresponding to periodic $x_{1}$, $x_{4}$ and 3:1 resonant orbits are labeled for reference.
@@ III.2 $x_{2}$ orbits
We adopt $M_{s}=0.3$ and $s=0.05$ for the spheroidal/bulge component, and $M_{s}=25$, $s=1.5$ for the halo component (J. A. Sellwood & A. Wilkinson, 1993). The pattern speed is set to be $\Omega_{\mathrm{b}}=2$.
@@
Figure 3: Same as Figure 2, but for the Ferrers bar potential. Chaotic are shown as scattered points, while quasi-periodic orbits appear as colored curves composed of discrete point. Distinct branches corresponding to periodic $x_{1}$, $x_{2}$, $x_{4}$ and 3:1 resonant orbits are labeled for reference.
@@ IV Test-particle orbits in the $\mathrm{CAM}$ framework
The simulation is set up by generating test particles with simple initial conditions as follows: i) The initial radius $R$ is drawn from an exponential distribution with scale length $r_{d}=2.2$, such that the probability density is proportional to $\exp(-R/r_{d})$. The initial azimuthal angle $\phi$ is uniformly distributed between 0 and $2\pi$. ii) The initial velocities in the radial and azimuthal directions are sampled from different Gaussian distributions. The mean radial velocity is set to zero, while the mean azimuthal velocity is equal to the rotation velocity of the corresponding potential model. For the Freeman bar and Ferrers bar potentials, the isotropic velocity dispersions are $0.4$. For the logarithmic bar potential, the isotropic velocity dispersion is set to be $0.5$. We integrate the orbits by using the python package Agama (E. Vasiliev, 2018).
@@ IV.2 Logarithmic bar potential
Orbits beyond the corotation radius also exhibit well-ordered distributions in our analysis. These orbits consistently display negative $\mathrm{CAM}$, as their rotation lags behind the bar pattern speed. Due to the averaging of the non-axisymmetric forces, they generally maintain nearly circular shapes. The outermost orbits possess the lowest $E_{\mathrm{J}}$ and form narrow, rosette-like annuli with small $|\mathrm{CAM}|$ values. In addition, we identify high-$E_{\mathrm{J}}$ orbits with $\mathrm{CAM}\sim 0$ in this region, primarily corresponding to unstable trajectories that would eventually escape the system under longer integration times.
@@ V.3 Application to $N$-body simulations
Here we briefly clarify how the method can be applied to particles in a realistic $N$-body simulation snapshot. For a given snapshot, one can extract the phase-space coordinates $(\mathbf{x}_{n},\mathbf{v}_{n})$ for particles, together with the frozen gravitational potential associated with the snapshot. To classify the orbital families represented by these particles, the following steps may be followed:
i) Calculate $\overline{L_{z}}$, $\sigma_{L_{z}}$ and $R_{\mathrm{rms}}$} for all particles in the snapshot.
@@ VI Conclusion
We thank James Binney for useful comments. The research presented here is partially supported by Office of Science and Technology, Shanghai Municipal Government under grant Nos 24DX1400100, ZJ2023-ZD-001; by the National Natural Science Foundation of China under grant Nos. 12533004, 12025302, 11773052; by China Manned Space Program with grant no. CMS-CSST-2025-A11; and by the “111” Project of the Ministry of Education of China under grant No. B20019. This work has made use of the Gravity Supercomputer at the Department of Astronomy, Shanghai Jiao Tong University.

### v2

@@ I Introduction
Our main motivation is to investigate the effectiveness of $\mathrm{CAM}$ for 2D orbital classification in rotating bar potentials. While QS21 focused primarily on the empirical and numerical properties of $\mathrm{CAM}$, we begin by investigating its analytical implications within the Freeman bar model, aiming to gain insight into the physical meaning of this proxy. Next, we evaluate the performance of $\mathrm{CAM}$ for orbital classification across three representative rotating bar potentials: the Freeman bar, the logarithmic bar, and the Ferrers bar. In addition to the $\overline{L_{z}}-\sigma_{L_{z}}$ plane studied by QS21, we introduce a new parameter plane defined by $\mathrm{CAM}$ and the root-mean-square radius (${R_{\mathrm{rms}}}=\sqrt{\overline{R^{2}}}$) [^1] Throughout this paper, we use $R_{\mathrm{rms}}$} to characterize the orbital size, but we do confirm that the mean absolute radius ($\overline{|R|}$) produces comparable results., where the average is taken over time along the orbit using a uniform time interval. We first explore the behavior of iso-$E_{\mathrm{J}}$ and periodic orbits within the $\mathrm{CAM}$ framework, finding that periodic orbits form well-defined branches in the $\mathrm{CAM}-{R_{\mathrm{rms}}}$ plane that effectively partition the space into domains corresponding to specific orbital families. We then extend the analysis to test-particle orbits rather than orbits constrained by fixed Jacobi energies, thereby demonstrating the applicability of this method to realistic $N$-body simulations. Our analyses show that different orbital families exhibit a well-ordered distribution in the $\mathrm{CAM}-{R_{\mathrm{rms}}}$ plane. These results highlight CAM as a robust and efficient tool for orbital classification.
@@ III Iso-$E_{\mathrm{J}}$ orbits and periodic orbits in the $\mathrm{CAM}$ framework
In the Freeman bar potential, the iso-$E_{\mathrm{J}}$ orbits in the $\overline{L_{z}}-\sigma_{L_{z}}$ plane are shown in the black curves in the left panel of Figure 1, forming the upper portion of an ellipse. The periodic $x_{1}$ and $x_{4}$ orbits in this model, however, exhibit zero angular momentum dispersion and constant $L_{z}$ as $X_{\alpha}$ or $X_{\beta}$ equals to 0 (see Equation 9), causing their $\mathrm{CAM}$ values to diverge. As a result, their properties cannot be meaningfully analyzed within the $\mathrm{CAM}$ framework. For this reason, we also consider two additional representative rotating bar potentials: the logarithmic bar and the Ferrers bar in J. A. Sellwood & A. Wilkinson (1993). The logarithmic bar model admits only the $x_{1}$ and $x_{4}$ orbital families, whereas the Ferrers bar potential also supports the $x_{2}$ orbital family. Orbits in these two potentials are integrated for a timespan of around $50$ dynamical timescales using the python package Agama, with a uniform time interval of $0.001$ dynamical timescales (E. Vasiliev, 2018). We have verified that extending or shortening the integration time by a factor of a few does not affect the main results. The periodic orbits are obtained by launching particles along the $x$-axis at fixed $E_{\mathrm{J}}$ and identifying those that return to their initial phase-space coordinates.
@@
In addition to the $\overline{L_{z}}-\sigma_{L_{z}}$ plane examined in the original CAM study QS21, we also explore alternative parameter spaces that directly incorporate the $\mathrm{CAM}$ value. Since CAM serves as a proxy for $I_{2}$ in addition to $E_{\mathrm{J}}$, the $\mathrm{CAM}-E_{\mathrm{J}}$ plane, resembling the integral-of-motion space, naturally provides a framework for orbital classification. However, because orbits inside and outside the corotation radius overlap in the $\mathrm{CAM}-E_{\mathrm{J}}$ plane, an issue discussed in detail in Section IV.2 later, we instead adopt the $\mathrm{CAM}-{R_{\mathrm{rms}}}$ plane, where $R_{\mathrm{rms}}$} is defined as the time-averaged root-mean-square radius. We find that this parameter space provides a clear separation between orbital families and serves as a critical diagnostic for orbital classification within the $\mathrm{CAM}$ framework.
@@ III.1 $x_{1}$ and $x_{4}$ orbits
$\Phi(x,y)=\frac{1}{2}v_{0}^{2}\ln(R_{c}^{2}+x^{2}+\frac{y^{2}}{b^{2}}).$ (22)
We take $v_{0}=1$, $R_{c}=0.1$ and $b=0.84$, with a pattern speed of $\Omega_{\mathrm{b}}=1$. The corotation radius and circular velocity at corotation are $r_{\rm CR}=0.99$ and $V_{\rm CR}=0.99$, respectively, owing to the nonzero value of $R_{c}$. Radii and velocities are therefore expressed in units of $1.01\ r_{\rm CR}$ and $1.01\ V_{\rm CR}$. This model only supports $x_{1}$ and $x_{4}$ orbital families, while $x_{2}$ and $x_{3}$ orbital families are absent.
@@
Figure 2: Orbital distributions of the logarithmic bar potential of Equation 22, shown in the ${\overline{L}_{z}}-\sigma_{L_{z}}$ plane (left) and the $\mathrm{CAM}-{R_{\mathrm{rms}}}$ plane (right), color-coded by $E_{\mathrm{J}}$. Representative periodic orbits, displayed with their morphological shapes, are superimposed on the distribution, while quasi-periodic orbits appear as colored curves composed of discrete points. Distinct branches corresponding to periodic $x_{1}$, $x_{4}$ and 3:1 resonant orbits are labeled for reference. The corotation radius is $1.01$, and the circular velocity and angular momentum at corotation are $1.01$ and $0.98$, respectively.
@@ III.2 $x_{2}$ orbits
We adopt $M_{s}=0.3$ and $s=0.05$ for the spheroidal/bulge component, and $M_{s}=25$, $s=1.5$ for the halo component (J. A. Sellwood & A. Wilkinson, 1993). The pattern speed is set to be $\Omega_{\mathrm{b}}=2$. The corotation radius and the circular velocity at corotation are $r_{\rm CR}=1.28$ and $V_{\rm CR}=2.56$, respectively. All radii and velocities are expressed in units of $0.78\ r_{\rm CR}$ and $0.39\ V_{\rm CR}$.
@@
Figure 3: Same as Figure 2, but for the Ferrers bar potential in Section III.2. Chaotic are shown as scattered points, while quasi-periodic orbits appear as colored curves composed of discrete point. Distinct branches corresponding to periodic $x_{1}$, $x_{2}$, $x_{4}$ and 3:1 resonant orbits are labeled for reference. The corotation radius is $1.28$, and the circular velocity and angular momentum at corotation are $2.56$ and $3.28$, respectively.
@@ IV Test-particle orbits in the $\mathrm{CAM}$ framework
The simulation is set up by generating test particles with simple initial conditions as follows: i) The initial radius $R$ is drawn from an exponential distribution with scale length $r_{d}=2.2$, such that the probability density is proportional to $\exp(-R/r_{d})$. We have verified that the specific choice of radial profile does not affect the final results. The initial azimuthal angle $\phi$ is uniformly distributed between 0 and $2\pi$. ii) The initial velocities in the radial and azimuthal directions are sampled from different Gaussian distributions. The mean radial velocity is set to zero, while the mean azimuthal velocity is equal to the rotation velocity of the corresponding potential model. The isotropic velocity dispersion is set to $0.4$ for the Freeman bar, $1$ for the Ferrers bar potentials, and $0.5$ for the logarithmic bar potential. These values are approximately half of the rotation velocity at corotation and are chosen to ensure a wide range of orbital behaviors, with the exception of the Freeman bar which does not possess a corotation radius. We further verify that varying the velocity dispersion by up to $50\%$ primarily affects the density of the orbital distribution, without altering its qualitative structure. We integrate the orbits by using the python package Agama (E. Vasiliev, 2018).
@@ IV.2 Logarithmic bar potential
Orbits beyond the corotation radius also exhibit well-ordered distributions in our analysis. These orbits consistently display negative $\mathrm{CAM}$, as their rotation lags behind the bar pattern speed. Due to the averaging of the non-axisymmetric forces, they generally maintain nearly circular shapes and form narrow, rosette-like annuli with large $|\mathrm{CAM}|$ values. The outermost orbits, which have large $R_{\mathrm{rms}}$}, correspond to the lowest values of $E_{\mathrm{J}}$ and are therefore more tightly bound. This behavior follows directly from the definition of the Jacobi energy:
$E_{\mathrm{J}}=\frac{1}{2}|\dot{\vec{r}}|^{2}+\Phi(\vec{r})-\frac{1}{2}\Omega_{\mathrm{b}}^{2}r^{2},$ (26)
where $\vec{r}$ and $\dot{\vec{r}}$ are defined in the corotation frame. For nearly circular orbits outside the corotation radius, the centrifugal term decreases more rapidly with increasing $R_{\mathrm{rms}}$} than the kinetic and gravitational potential terms, leading to lower $E_{\mathrm{J}}$ for larger $R_{\mathrm{rms}}$}.
In addition, we identify high-$E_{\mathrm{J}}$ orbits with $\mathrm{CAM}\sim 0$ in this region, primarily corresponding to unstable trajectories that would eventually escape the system under longer integration times.
@@ V.3 Application to $N$-body simulations
Here we briefly clarify how the method can be applied to particles in a realistic $N$-body simulation snapshot. For a given snapshot, one can extract the phase-space coordinates $(\mathbf{x}_{n},\mathbf{v}_{n})$ for particles, together with the frozen gravitational potential associated with the snapshot. To classify the orbital families represented by these particles, the following steps can be adopted:
i) Integrate snapshot particles in the corresponding frozen potential. The quantities $\overline{L_{z}}$, $\sigma_{L_{z}}$, and $R_{\mathrm{rms}}$} can be computed by time-averaging along the resulting orbits, and $\mathrm{CAM}$ is defined as the ratio $\overline{L_{z}}/\sigma_{L_{z}}$.
@@ VI Conclusion
We thank James Binney for useful comments. The research presented here: is partially supported by the National Natural Science Foundation of China under grant Nos. 12533004, 12025302, 11773052; by China Manned Space Program with grant no. CMS-CSST-2025-A11; by the “111” Project of the Ministry of Education of China under grant No. B20019; and by Office of Science and Technology, Shanghai Municipal Government with grant Nos. 24DX1400100 and ZJ2023-ZD-001. This work has made use of the Gravity Supercomputer at the Department of Astronomy, Shanghai Jiao Tong University.